OpenStudy (anonymous):

Simplify completely the quantity 10 times x to the 6th power times y to the third power plus 20 times x to the third power times y to the 2nd power all over 5 times x to the third power times y

OpenStudy (anonymous):

@welshfella

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

can you help me @pandaluvs

OpenStudy (anonymous):

I'm sorry i suck at this stuff =)

OpenStudy (anonymous):

do u know anyone that can help me

OpenStudy (anonymous):

please

OpenStudy (anonymous):

@MarshallEatsFood @Great_And_Powerful @ribhu @Love_Ranaa

OpenStudy (love_ranaa):

@k_lynn ??

OpenStudy (anonymous):

@StudyGurl14

OpenStudy (anonymous):

I need help with 5 questions like this anyone please omgg

OpenStudy (studygurl14):

Like this? \(\Large \frac{10x^6y^3+20x^3y^2}{5x^3y}\) 10 times x to the 6th power times y to the third power plus 20 times x to the third power times y to the 2nd power all over 5 times x to the third power times y

OpenStudy (anonymous):

yes

OpenStudy (studygurl14):

First step. Factor of \(\large 5x^3y\) from the numerator so you can eliminate the denominator. \(\huge\frac{10x^6y^3+20x^3y^2}{5x^3y}\rightarrow\frac{5x^3y(2x^3y^2+4y)}{5x^3y}\rightarrow ?\) Can you do the rest?

OpenStudy (anonymous):

no i really dnt get this stuff

OpenStudy (anonymous):

nd i also need help with four more

OpenStudy (studygurl14):

the \(5x^3y\) in the numerator and the denominator can cancel out because... \(\Large \frac{5x^3y}{5x^3y}=\large1\) So...you can simplify the expression into... \(\large 2x^3y^2+4y\)

OpenStudy (studygurl14):

That's it, unless they also want you to factor out the 2y

OpenStudy (anonymous):

thank you

OpenStudy (studygurl14):

You're welcome.